OpenStudy (anonymous):

Which of the following vectors are orthogonal to (2,1)? **choose ALL that apply!!!** A. (-3,6) B. (1,-2) C. (-2,-3) D. (1,2) **a bit confused on this.. not too sure how to figure out if they are orthogonal :/ @terenzreignz :)

terenzreignz (terenzreignz):

Dot product I just mentioned that earlier ... YOU just mentioned that earlier...

OpenStudy (anonymous):

\[v* u=0\]

OpenStudy (anonymous):

haha oops :P its not that ginormous equation thingy earlier right? :/

OpenStudy (anonymous):

*from earlier

terenzreignz (terenzreignz):

No, that was cross product, thank heavens we don't have to deal with that right now... dot product of a pair of 2d vectors is far simpler... <a , b> dot <c , d> = ac + bd

OpenStudy (anonymous):

haha ohh okay :P yay! sooo 2(-3) + 1(6) = -6 +6 = 0 so A is a solution?

OpenStudy (anonymous):

\[(x_1,y_1)* (x_2,y_2)=x_1x_2+y_1y_2\]

terenzreignz (terenzreignz):

So,for instance, what's the dot product of <2 , 1> with <-3 , 6> ?

OpenStudy (anonymous):

2(1) + 1(-2) = 2 + (-2) = 2-2 = 0 so B is a solution also? :/

OpenStudy (anonymous):

ermm -6 + 6 =0 ?

terenzreignz (terenzreignz):

Yup. As long as their dot product is zero, they are orthogonal (I'm pretty sure I already mentioned that :D )

OpenStudy (anonymous):

ahh i see.. kk so we have A and B as solutions! lemme check the other two! :D

OpenStudy (anonymous):

2(-2) + 1(-3) = -4 -3 = -7 so C is NOT an answer :P

OpenStudy (anonymous):

2(1) + 1(2) = 2 + 2 =4 so again, nope! so just A and B then! :)

terenzreignz (terenzreignz):

GREAT :D

OpenStudy (anonymous):

thanks!! :D

OpenStudy (anonymous):

thanks!! :D